OpenStudy (anonymous):

Factor the trinomial completely t^2 + (2/9)t + (1/81)

OpenStudy (trojanpoem):

(t + 1/9)^2

OpenStudy (anonymous):

how did you do that?

OpenStudy (trojanpoem):

See the middle term 2 * 1/9 * t the first term is just the t^2 the last term is the (1/9)^2 that's on the form a^2 + 2ab + b^2 = (a+b)^2

OpenStudy (anonymous):

where does the 2 go in 2 * 1/9 * t? how did you plug it into that formula?

OpenStudy (trojanpoem):

We have 2ab = 2/9 t ab = 1/9 t a = t b = 1/9 The same as the product we found so we can apply the formula a^2 + 2ab + b^2 = (a+b)^2 = (t + 1/9)^2

OpenStudy (anonymous):

thank you!

OpenStudy (trojanpoem):

Any time.

OpenStudy (anonymous):

k